Question 12

Alpha Mining Company recognizes $2 of depletion for each ton of ore mined. During 2016, 850,000 tons of ore were mined and 725,000 tons were sold. Using this information, how much should be recorded on June 30 for the following accounts:

Accumulated Depletion

Inventory

Cost of Goods Sold
Answer #1

Accumulated depreciation= no.of tons extracted*depletion rate 850000*$2=$1700000

Inventory no.of units extracted-no. Of tons sold*depletion rate= ( 850000-725000)*2=125000*2=$250000

Cost of good sold= no. of tons sold* depletion rate= 725000*2=$1450000

If any doubt please comment